The U.S. and Mexico both produce vehicles and tons of plastic, which are sold for the same price in both countries. Suppose that with one unit of labor and one unit of capital, the U.S. can produce either 20 vehicles or 40 tons of plastic and Mexico can produce either 10 vehicles or 15 tons of plastic. Which country has a comparative advantage producing plastic?
Answer #1

Opportunity cost of producing plastic in US = 20/40 = 0.5 vehicles

Opportunity cost of producing plastic in Mexico = 10/15 = 0.66 vehicles

So, US has a comparative advantage in the production of plastic